OpenStudy (anonymous):

Name the slope y=4x-6

OpenStudy (acxbox22):

y=mx+b where m is slope and b is y intercept using this form the slope is 6

OpenStudy (anonymous):

no it's not

OpenStudy (anonymous):

right terms wrong answer

OpenStudy (dangerousjesse):

The slope is four, actually.

OpenStudy (anonymous):

y=mx+b m=slope b=y-intercept what number is substituted for m?

OpenStudy (acxbox22):

sorry it is 4... was in a rush

OpenStudy (anonymous):

compare it with y=mx+c here m is slope/////////